How much alcohol must be added to 480 grams of hand sanitizer that is 24% alcohol to
make it a hand sanitizer that is 40% alcohol? Correct your answer to the nearest whole
number
1
Expert's answer
2020-06-10T19:59:00-0400

Solution.

Let it is added x gram of 100% alcohol. Find the weight of alcohol, which must be added:

"x*1+480*0.24=(x+480)*0.4;\\
x+115.2=0.4x+192;\\
0.6*x=76.8;\\
x=128.\\"

Answer: 128 gram.